# DD-Ranking: Rethinking the Evaluation of Dataset Distillation

### v2

@@ 1 Introduction
With the rapid advancement of deep learning, training increasingly complex and more complex models on large scale datasets has become a standard paradigm, achieving remarkable performance in various fields, such as computer vision [20, 9] and natural language processing [8, 1]. However, this process often incurs substantial computational and storage demands, significantly hindering deployment across diverse scenarios. Dataset distillation (DD) [54], as a recent promising solution for dataset compression, offers novel insights to address these challenges. In recent years, diverse training pipelines [21, 7, 64, 15, 59] and optimization objectives [67, 63, 2] have been proposed, driving rapid advancement in the field of dataset distillation.
To further enhance the testing accuracy of models trained on synthetic datasets during the post-evaluation phase, recent studies have incorporated general performance boosting techniques (e.g., soft labels) into the evaluation process. Some methods jointly optimize the generated images and their corresponding unique soft labels [18, 31], while decoupled dataset distillation methods [59, 41, 46, 48, 37] utilize epoch-wise soft labels provided by pre-trained teacher models during post-evaluation phase. Although these works successfully demonstrate that soft labels significantly improve testing accuracy of the validation models, their soft label implementation strategies differ substantially, and performance comparisons with prior methods often fail to account for gains attributable to soft labels.
Furthermore, Subsequent studies frequently employ more intensive data augmentation, superior optimizers, and refined training hyper-parameters [43, 4] during evaluation to maximize model performance, with even randomly sampled images achieving superior results under better post-evaluation settings. This practice conflates genuine improvements in dataset quality with performance variations caused by inconsistent evaluation settings, severely impeding progress in dataset distillation and directing subsequent improvements toward suboptimal directions. Based on the aforementioned discussion, we must emphasize that in the growing field of dataset distillation, relying solely on the testing accuracy of validation model as the exclusive criterion for assessing the quality of synthetic datasets exhibits significant unreliability and unfairness when applied across varying settings.
@@ 2.2 Soft Labels
Soft labels have been extensively employed in various domains, particularly in knowledge distillation tasks. Unlike hard labels, which assign discrete categorical values, soft labels represent probability distributions across class categories. These distributions are typically derived from the output logits of pretrained models. Recently, applying soft labels has emerged as a popular approach in evaluating dataset distillation methods. In this framework, each distilled image is associated with one or multiple soft labels generated by a pretrained teacher model. For instance, DATM [18] concurrently optimizes synthetic data and corresponding soft labels during bi-level optimization procedures. SRe2L [59] employs a teacher model to generate multiple soft labels per data sample at test time. Consequently, the training objective for an agent model on the distilled dataset becomes minimizing the loss (e.g., Kullback–Leibler divergence) between its output logits and these soft labels. Due to this knowledge distillation paradigm, evaluation metrics with soft labels consistently demonstrate substantially higher performance, as illustrated in Figure 1.
@@
Improvements originate from knowledge distillation [37], instead of synthetic data.
@@ 3.2 Label-Robust Score
The initial goal of dataset distillation is to synthesize a small number of data points that do not need to come from the correct data distribution, but will, when given to the learning algorithm as training data, approximate the model trained on the original data [54]. Given that almost all existing classification datasets use hard label annotation, we think it is crucial for DD methods to maintain good performance with hard labels. To this end, we propose the hard label recovery (HLR). Specifically, for both hard-label-based and soft-label-based methods, we evaluate the test accuracy of the distilled data and that of the original dataset with hard labeling, denoted as $\text{acc}_{\text{syn-hard}}$ and $\text{acc}_{\text{real-hard}}$, respectively. The hard label recovery is computed by taking the difference:
@@
We assign a negative mark to HLR so that both parts of the sum are positively correlated with the performance. The raw range of $\alpha$ is between $[-1,1]$, so we normalize LRS to the range $[0,1]$ by letting $\text{LRS}=100\%\times(e^{\alpha}-e^{-1})/(e-e^{-1})$. A higher LRS indicates that the distilled dataset of the corresponding method is more robust to the label representation and has richer information.
@@ 4.1 Evaluation Settings
We evaluate a wide range of representative works in dataset distillation. For hard-label methods, we evaluate DC [67], DSA [62], MTT [2], DM [63], and DataDAM [39]. For soft-label methods, we evaluate SRe2L [59], DATM [18], EDF [50], DWA [13], RDED [48], CDA [58], EDC [43], and G-VBSM [42]. In the case where the method provides its distilled data, we adopt it directly. In the case where the distilled data is absent, we strictly follow their implementation provided in both the paper and code repo to replicate their results.
@@
We report DD-Ranking benchmarking results on the four existing datasets: CIFAR-10 [22], CIFAR-100 [22], TinyImageNet [24], and ImageNet1K [38]. The resolution of images in CIFAR-10 and CIFAR-100 is $32\times 32$. The resolution of images in TinyImageNet is $64\times 64$. The resolution of images in ImageNet1K is $224\times 224$. We only report ARS results on ImageNet1K due to space limit. More results can be found in our leaderboard.
@@
For each baseline method, we use the model architecture reported in the paper for evaluation. This includes ConvNet of depth 3 and 4 with instance normalization, ConvNet of depth 3 and 4 with batch normalization, and ResNet-18 [19]. Additionally, to validate the robustness of DD-Ranking on different model architectures, we incorporate AlexNet [23], ResNet-50, VGG-11 [45], Swin-Transformer-tiny [30], and Vision-Transformer-base [10].
@@
The evaluation is performed 5 times with different random seeds. We report the mean value in the following tables. When computing the accuracy under hard labels, we perform the hyperparameter search for the learning rate and report the best one. When computing the accuracy under soft labels, we regard the learning provided by each method as the optimal learning rate by default, and the learning rate search is performed for random selection.
@@ 4.4 Analysis
Various dataset distillation methods employ different agent models for distillation architecture or evaluation of synthetic data. Despite variations in raw test accuracy across model architectures, the metric used to evaluate dataset distillation performance should remain consistent, with minimal fluctuation in metric values. As shown in Figure 2, the Learning Rate Schedule (LRS) results for six methods across different settings, each tested with four distinct model architectures, demonstrate high consistency. This consistency validates the robustness of our benchmark across different model architectures.
@@
In decoupled dataset distillation [59, 41, 46, 48], epoch-wise soft labels constitute a crucial component of the synthetic dataset. Recent studies [41, 43, 4] have explored improving test accuracy by leveraging stronger teacher models to provide soft labels without altering the synthetic data itself. However, the validity of this technique remains insufficiently investigated. As shown in Figure 3, whether through the use of different teacher models or advanced hybrid soft label strategies by fusing soft labels generated by multiple teachers, our proposed LRS consistently exhibits strong robustness, thereby validating its reliability across diverse soft label settings.
@@ 5 Related Works
Hard-label-based DD methods assign categorical labels to synthetic samples, the same as the labels of the real dataset. Matching-based methods are known as representative hard-label-based methods. i) Gradient matching: Synthetic data is optimized such that the gradients they induce on a neural network closely match those from real data. Following the pioneering work of Dataset Condensation (DC) [67], various works have improved gradient matching, such as DSA [62], DCC [25], and LCMat [44]. ii) Trajectory matching: Synthetic data are optimized by aligning the training dynamics of models trained on synthetic data with those trained on real data. MTT [2] first introduced this approach, where synthetic data is optimized by aligning the training dynamics of models trained on synthetic data with those trained on real data. Building on this, various methods such as TESLA [6], FTD [11], and ATT [28] further enhance trajectory matching by improving memory efficiency and reducing trajectory errors. iii) Feature matching is an alternative to gradient or trajectory-based distillation, where synthetic data is optimized to induce similar internal representations as real data. Represented by CAFE [51], DM [66], and DataDAM [39], this approach offers a lightweight framework with comparable performances, especially on large IPC settings.
@@
. DD methods using soft labels employ knowledge distillation during evaluation. Each synthetic sample is assigned to one or multiple soft labels generated by a pretrained teacher model. Among matching-based methods, DATM [18], PAD [27], and EDF [50] optimize the soft labels jointly with synthetic data during trajectory matching. Recently, decoupled methods have demonstrated strong scalability on large datasets such as ImageNet1K by decoupling the bi-level optimization. SRe2L [59] first proposed a three-stage "squeeze, recover, and relabel" paradigm. During the relabel stage, soft labels are generated and saved for each synthetic sample. Following this approach, CDA [58], DWA [48], EDC [43], and G-VBSM [42] further improve the performance from both data and soft label perspectives. RDED synthesizes condensed data by concatenating core image patches. D4M employs diffusion models to generate high-quality synthetic images.
@@
A notable challenge for dataset distillation is the lack of comprehensive benchmarks. DC-Bench [5] is the first large-scale standardized benchmark for dataset condensation methods in general. It provides a comprehensive evaluation for several dataset distillation methods and coreset selection methods. Comp-DD is proposed in EDF [50] targeting dataset distillation in complex scenarios. It extracts new subsets from ImageNet1K based on the complexity metric. However, both benchmarks no longer satisfy the need for fair evaluation of dataset distillation methods under the soft label trend. Therefore, DD-Ranking is proposed to solve this problem.
@@ 6 Conclusion and Future Work
We propose DD-Ranking, a new benchmark that provides a fair and comprehensive evaluation for dataset distillation. DD-Ranking is well motivated by the unfairness originated from inconsistent training settings of existing DD evaluation, especially the use of soft labels and data augmentation. To this end, DD-Ranking introduces both label robust score and augmentation robust score to disentangle the effect of knowledge distillation via soft labeling and data augmentation, and ultimately reveal the true informativeness of distilled datasets. Hopefully, DD-Ranking can facilitate the development of dataset distillation towards improving data quality instead of accuracy. DD-Ranking is already open-source as a PyPI package with detailed documentation. One potential limitation of the current DD-Ranking is that we only support methods for image classification dataset distillation. We are aware that several works [72, 55] have extended dataset distillation to other tasks and modalities. In the future, we will constantly integrate more baseline methods into our benchmark and extend DD-Ranking to other modalities.
@@ Appendix B Additional Experiment Results
Results from Table 3 to Table 6 are computed by letting $\lambda=0.5$. By default, we treat the hard label recovery and improvement over random equally important. In Table 8 to Table 19, we report the LRS results under different $\lambda$. A larger $\lambda$ gives higher priority to IOR, and a smaller $\lambda$ focuses more on HLR. We encourage future newly proposed DD methods to enhance both HLR and IOR.

### v3

@@ 1 Introduction
With the rapid advancement of deep learning, training increasingly complex and more complex models on large scale datasets has become a standard paradigm, achieving remarkable performance in various fields, such as computer vision he2016deep ; dosovitskiy2020image and natural language processing devlin2018bert ; brown2020language . However, this process often incurs substantial computational and storage demands, significantly hindering deployment across diverse scenarios. Dataset distillation (DD) wang2018dataset , as a recent promising solution for dataset compression, offers novel insights to address these challenges. In recent years, diverse training pipelines kim2022dataset ; NEURIPS2022_de3d2bb6 ; zhao2022synthesizing ; feng2024embarrassingly ; yin2024squeeze and optimization objectives zhao2021dataset ; zhao2021datasetdm ; cazenavette2022dataset have been proposed, driving rapid advancement in the field of dataset distillation.
To further enhance the testing accuracy of models trained on synthetic datasets during the post-evaluation phase, recent studies have incorporated general performance boosting techniques (e.g., soft labels) into the evaluation process. Some methods jointly optimize the generated images and their corresponding unique soft labels guo2023towards ; loo2022efficient , while decoupled dataset distillation methods yin2024squeeze ; shao2024generalized ; su2024d4m ; sun2024diversity ; qin2024label utilize epoch-wise soft labels provided by pre-trained teacher models during post-evaluation phase. Although these works successfully demonstrate that soft labels significantly improve testing accuracy of the validation models, their soft label implementation strategies differ substantially, and performance comparisons with prior methods often fail to account for gains attributable to soft labels.
Furthermore, subsequent studies frequently employ more intensive data augmentation, superior optimizers, and refined training hyper-parameters shao2024elucidating ; cui2025dataset during evaluation to maximize model performance, with even randomly sampled images achieving superior results under better post-evaluation settings. This practice conflates genuine improvements in dataset quality with performance variations caused by inconsistent evaluation settings, severely impeding progress in dataset distillation and directing subsequent improvements toward suboptimal directions. Based on the aforementioned discussion, we must emphasize that in the growing field of dataset distillation, relying solely on the testing accuracy of validation model as the exclusive criterion for assessing the quality of synthetic datasets exhibits significant unreliability and unfairness when applied across varying settings.
@@ 2.2 Soft Labels
Soft labels have been extensively employed in various domains, particularly in knowledge distillation tasks. Unlike hard labels, which assign discrete categorical values, soft labels represent probability distributions across class categories. These distributions are typically derived from the output logits of pretrained models. Recently, applying soft labels has emerged as a popular approach in evaluating dataset distillation methods. In this framework, each distilled image is associated with one or multiple soft labels generated by a pretrained teacher model. For instance, DATM (guo2023towards, ) concurrently optimizes synthetic data and corresponding soft labels during bi-level optimization procedures. SRe2L (yin2024squeeze, ) employs a teacher model to generate multiple soft labels per data sample at test time. Consequently, the training objective for an agent model on the distilled dataset becomes minimizing the loss (e.g., Kullback–Leibler divergence) between its output logits and these soft labels. Due to this knowledge distillation paradigm, evaluation metrics with soft labels consistently demonstrate substantially higher performance, as illustrated in Figure 1.
@@
Improvements originate from knowledge distillation qin2024label , instead of synthetic data.
@@ 3.2 Label-Robust Score
The initial goal of dataset distillation is to synthesize a small number of data points that do not need to come from the correct data distribution, but will, when given to the learning algorithm as training data, approximate the model trained on the original data (wang2018dataset, ). Given that almost all existing classification datasets use hard label annotation, we think it is crucial for DD methods to maintain good performance with hard labels. To this end, we propose the hard label recovery (HLR). Specifically, for both hard-label-based and soft-label-based methods, we evaluate the test accuracy of the distilled data and that of the original dataset with hard labeling, denoted as $\text{acc}_{\text{syn-hard}}$ and $\text{acc}_{\text{real-hard}}$, respectively. The hard label recovery is computed by taking the difference:
@@
We assign a negative mark to HLR so that both parts of the sum are positively correlated with the performance. The raw range of $\alpha$ is between [1,1][-1,1], so we normalize LRS to the range $[0,1]$ by letting $\text{LRS}=100\%\times(e^{\alpha}-e^{-1})/(e-e^{-1})$. A higher LRS indicates that the distilled dataset of the corresponding method is more robust to the label representation and has richer information.
@@ 4.1 Evaluation Settings
We evaluate a wide range of representative works in dataset distillation. For hard-label methods, we evaluate DC (zhao2021dataset, ), DSA (zhao2021dsa, ), MTT (cazenavette2022dataset, ), DM (zhao2021datasetdm, ), and DataDAM (sajedi2023datadam, ). For soft-label methods, we evaluate SRe2L (yin2024squeeze, ), DATM (guo2023towards, ), EDF (wang2025emphasizingdiscriminativefeaturesdataset, ), DWA (du2024diversitydriven, ), RDED (sun2024diversity, ), CDA (yin2024dataset, ), EDC (shao2024elucidating, ), and G-VBSM (shao2024generalizedlargescaledatacondensatio, ). In the case where the method provides its distilled data, we adopt it directly. In the case where the distilled data is absent, we strictly follow their implementation provided in both the paper and code repo to replicate their results.
@@
We report DD-Ranking benchmarking results on the four existing datasets: CIFAR-10 (Krizhevsky2009LearningML, ), CIFAR-100 (Krizhevsky2009LearningML, ), TinyImageNet (Le2015TinyIV, ), and ImageNet1K (russakovsky2015imagenetlargescalevisual, ). The resolution of images in CIFAR-10 and CIFAR-100 is $32\times 32$. The resolution of images in TinyImageNet is $64\times 64$. The resolution of images in ImageNet1K is $224\times 224$. We only report ARS results on ImageNet1K due to space limit. More results can be found in our leaderboard.
@@
For each baseline method, we use the model architecture reported in the paper for evaluation. This includes ConvNet of depth 3 and 4 with instance normalization, ConvNet of depth 3 and 4 with batch normalization, and ResNet-18 (He2015DeepRL, ). Additionally, to validate the robustness of DD-Ranking on different model architectures, we incorporate AlexNet (AlexNet, ), ResNet-50, VGG-11 (VGG11, ), Swin-Transformer-tiny (Swin-T, ), and Vision-Transformer-base (ViT, ).
@@
The evaluation is performed 5 times with different random seeds. We report the mean value in the following tables. Standard deviations are reported in the appendix. When computing the accuracy under hard labels, we perform the hyperparameter search for the learning rate and report the best one. When computing the accuracy under soft labels, we regard the learning provided by each method as the optimal learning rate by default, and the learning rate search is performed for random selection.
@@ 4.4 Analysis
Cross-architecture evaluation is an important experiment for dataset distillation methods. Specifically, different models architectures are used to evaluate the synthetic data. Despite variations in raw test accuracy across model architectures, the metric used to evaluate dataset distillation performance should remain consistent, with minimal fluctuation in metric values. As shown in Figure 2, the LRS results for six methods across different settings, each tested with four distinct model architectures, demonstrate high consistency. This consistency validates the robustness of our benchmark across different model architectures.
@@
In decoupled dataset distillation yin2024squeeze ; shao2024generalized ; su2024d4m ; sun2024diversity , epoch-wise soft labels constitute a crucial component of the synthetic dataset. Recent studies shao2024generalized ; shao2024elucidating ; cui2025dataset have explored improving test accuracy by leveraging stronger teacher models to provide soft labels without altering the synthetic data itself. However, the validity of this technique remains insufficiently investigated. As shown in Figure 3, whether through the use of different teacher models or advanced hybrid soft label strategies by fusing soft labels generated by multiple teachers, our proposed LRS consistently exhibits strong robustness, thereby validating its reliability across diverse soft label settings.
@@ 5 Related Works
Hard-label-based DD methods assign categorical labels to synthetic samples, the same as the labels of the real dataset. Matching-based methods are known as representative hard-label-based methods. i) Gradient matching: Synthetic data is optimized such that the gradients they induce on a neural network closely match those from real data. Following the pioneering work of Dataset Condensation (DC) zhao2021dataset , various works have improved gradient matching, such as DSA (zhao2021dsa, ), DCC (DCC, ), and LCMat (LCMat, ). ii) Trajectory matching: Synthetic data are optimized by aligning the training dynamics of models trained on synthetic data with those trained on real data. MTT cazenavette2022dataset first introduced this approach, where synthetic data is optimized by aligning the training dynamics of models trained on synthetic data with those trained on real data. Building on this, various methods such as TESLA cui2023scaling , FTD (du2023minimizing, ), and ATT (ATT, ) further enhance trajectory matching by improving memory efficiency and reducing trajectory errors. iii) Feature matching is an alternative to gradient or trajectory-based distillation, where synthetic data is optimized to induce similar internal representations as real data. Represented by CAFE wang2022cafe , DM zhao2023dataset , and DataDAM sajedi2023datadam , this approach offers a lightweight framework with comparable performances, especially on large IPC settings.
@@
. DD methods using soft labels employ knowledge distillation during evaluation. Each synthetic sample is assigned to one or multiple soft labels generated by a pretrained teacher model. Among matching-based methods, DATM (guo2023towards, ), PAD (li2024prioritizealignmentdatasetdistillation, ), and EDF (wang2025emphasizingdiscriminativefeaturesdataset, ) optimize the soft labels jointly with synthetic data during trajectory matching. Recently, decoupled methods have demonstrated strong scalability on large datasets such as ImageNet1K by decoupling the bi-level optimization. SRe2L (yin2024squeeze, ) first proposed a three-stage "squeeze, recover, and relabel" paradigm. During the relabel stage, soft labels are generated and saved for each synthetic sample. Following this approach, CDA (yin2024dataset, ), DWA (sun2024diversity, ), EDC (shao2024elucidating, ), and G-VBSM (shao2024generalizedlargescaledatacondensatio, ) further improve the performance from both data and soft label perspectives. RDED synthesizes condensed data by concatenating core image patches. D4M employs diffusion models to generate high-quality synthetic images.
@@
A notable challenge for dataset distillation is the lack of comprehensive benchmarks. DC-Bench (cui2022dc, ) is the first large-scale standardized benchmark for dataset condensation methods in general. It provides a comprehensive evaluation for several dataset distillation methods and coreset selection methods. Comp-DD is proposed in EDF (wang2025emphasizingdiscriminativefeaturesdataset, ) targeting dataset distillation in complex scenarios. It extracts new subsets from ImageNet1K based on the complexity metric. However, both benchmarks no longer satisfy the need for fair evaluation of dataset distillation methods under the soft label trend. Therefore, DD-Ranking is proposed to solve this problem.
@@ 6 Conclusion and Future Work
We propose DD-Ranking, a new benchmark that provides a fair and comprehensive evaluation for dataset distillation. DD-Ranking is well motivated by the unfairness originated from inconsistent training settings of existing DD evaluation, especially the use of soft labels and data augmentation. To this end, DD-Ranking introduces both label robust score and augmentation robust score to disentangle the effect of knowledge distillation via soft labeling and data augmentation, and ultimately reveal the true informativeness of distilled datasets. Hopefully, DD-Ranking can facilitate the development of dataset distillation towards improving data quality instead of accuracy. DD-Ranking is already open-source as a PyPI package with detailed documentation. One potential limitation of the current DD-Ranking is that we only support methods for image classification dataset distillation. We are aware that several works (zhou2023datasetquantization, ; wu2024visionlanguagedatasetdistillation, ) have extended dataset distillation to other tasks and modalities. In the future, we will constantly integrate more baseline methods into our benchmark and extend DD-Ranking to other modalities.
@@ Appendix B Additional Experiment Results
Results from Table 3 to Table 6 are computed by letting $\lambda=0.5$. By default, we treat the hard label recovery and improvement over random equally important. In Table 8 to Table 19, we report the LRS results under different $\lambda$. A larger $\lambda$ gives higher priority to IOR, and a smaller $\lambda$ focuses more on HLR. We encourage future newly proposed DD methods to enhance both HLR and IOR. From Table 20(c) to Table 20(d), we provide the standard deviations for all benchmark results computed under 5 runs with different random seeds.
@@
ipc 1 10 50
[…]
DC 0.2 0.2 0.3 0.3 0.2 0.3 0.2 0.4 0.3
DSA 0.3 0.3 0.2 0.4 0.2 0.2 0.5 0.4 0.4
MTT 0.5 0.7 0.6 0.6 0.8 0.8 0.5 0.4 0.5
DM 0.7 0.6 0.5 0.8 0.9 1.0 0.7 0.7 0.7
DATADAM 0.8 0.5 0.6 0.6 0.6 0.5 0.7 0.5 0.6
DATM 0.7 0.4 0.8 0.5 0.7 0.6 0.3 0.7 0.5
SRe2L 0.5 0.6 0.6 0.8 0.8 0.6 0.5 0.8 0.7
RDED 0.7 0.9 0.7 0.8 0.7 0.8 0.9 1.2 0.9
D4M 0.8 0.8 0.6 0.7 0.9 0.9 0.8 1.0 0.9
(a) Standard deviations of LRS results on CIFAR-10.
ipc 1 10 50
[…]
DC 0.4 0.3 0.3 0.3 0.5 0.4 0.2 0.6 0.7
DSA 0.5 0.5 0.5 0.4 0.4 0.3 0.4 0.5 0.4
MTT 0.5 0.6 0.5 0.7 0.7 0.6 0.5 0.6 0.6
DM 0.6 0.8 0.7 0.9 0.9 0.9 0.7 0.7 0.5
DATADAM 0.6 0.7 0.8 0.5 0.8 0.7 0.7 0.6 0.6
DATM 0.7 0.5 0.6 0.6 0.6 0.6 0.5 0.8 0.7
SRe2L 1.1 0.9 0.9 0.8 0.7 0.7 0.5 0.9 0.7
RDED 0.7 1.0 0.8 0.6 0.9 0.6 0.8 1.1 0.9
D4M 0.5 0.6 0.4 0.7 0.8 0.6 0.8 0.9 0.9
(b) Standard deviations of LRS results on CIFAR-100.
ipc 1 10 50
[…]
DC 0.4 0.3 0.3 0.2 0.5 0.5 0.4 0.6 0.4
DSA 0.5 0.4 0.6 0.3 0.7 0.4 0.5 0.5 0.4
MTT 0.5 0.6 0.6 0.4 0.7 0.6 0.3 0.8 0.6
DM 0.3 0.7 0.5 0.8 0.9 0.8 0.5 0.8 0.7
DATM 0.5 0.4 0.4 0.3 0.5 0.6 0.4 0.9 0.6
SRe2L 0.7 0.7 0.7 0.5 0.4 0.5 0.4 0.8 0.8
RDED 0.6 0.8 0.7 0.5 0.9 0.7 0.7 1.0 0.9
D4M 0.6 0.7 0.6 0.7 0.8 0.6 0.9 1.1 0.8
(c) Standard deviations of LRS results on TinyImageNet.
ipc 1 10 50
[…]
SRe2L 0.6 1.1 0.8 0.5 0.8 0.9 0.7 1.0 0.7
RDED 0.5 0.7 0.6 0.4 0.9 0.8 0.7 0.7 0.7
D4M 0.8 0.6 0.6 0.5 1.1 0.9 0.6 0.8 0.5
DWA 0.8 1.2 1.1 0.9 1.1 1.0 0.7 0.8 0.9
CDA 0.6 0.6 0.8 0.7 0.8 0.7 0.4 0.8 0.6
EDC 0.3 0.8 0.4 0.5 0.4 0.5 0.5 1.1 0.9
G-VBSM 0.6 1.2 1.0 0.5 1.3 0.9 0.6 0.9 0.8
(d) Standard deviations of LRS results on ImageNet-1K.